### v2

# Beyond Pixels: Exploring DOM Downsampling for LLM-Based Web Agents

Beyond Pixels: Exploring DOM Downsampling for LLM-Based Web Agents
Thassilo M. Schiepanski
1 Introduction
Ever since the web’s inception, autonomous agents that browse the web have been a desire. Web agents are expected to act on UIs (user interfaces), rather than APIs (application programming interfaces). Not least to integrate a human-in-the-loop, but for a human-centric web. Conventional agents search formal models of a problem domain Browne et al. (2012). Think of a chessboard, or an apartment worked by a vacuum robot. The human-centric web, a plethora of heterogeneous web application UIs, scales beyond a manageable domain model.
Frontier LLMs recently gained ability to respond according to schema Pokrass (2024). This guarantee declares LLMs instantaneous backends for virtually any application. For instance, the state of a chess game could as well be modelled with natural language. The idea of LLM backends only recently enabled serviceable web agents, spawning the latest trend in agentic AI OpenAI (2025d, a); Müller and Žunič (2024); Anthropic (2024). In particular, generalist agents that are not constrained by application domain He et al. (2024); Zheng et al. (2024).
1.1 Problem
The key problem of LLM-based web agents – model and browser exist for reuse – is providing valuable model context OpenAI (a); Anthropic . At the core, this is, serialisation of web application state – herein referred to as snapshot. State-of-the-art web agents are premised on grounded GUI snapshots (screenshots) Zheng et al. (2024); OpenAI (2025d); Anthropic (2024); Müller and Žunič (2024); Zheng et al. . Evidently, a screenshot resembles how humans perceive a web application at a given point in time. Behind LLM APIs, however, image pre-processing irreversibly affects dimensions, which renders precise targeting of elements impossible OpenAI (2025c). Grounding means adding visual cues, commonly bounding boxes with symbolic identifiers, to allow targeting via identifier He et al. (2024). While vision capabilities are limited OpenAI (2025c), frontier LLM’s were supported strong abilities to interpret HTML, and even an inherent UI OpenAI (2025c); Gur et al. (2022). The DOM (document object model) – a web application’s runtime state model Marini (2002); MDN Contributors (a) – serialises to HTML MDN Contributors (c). That said, DOM snapshots impose an alternative to GUI snapshots.
[…]
1.2 Contributions
In this work, we propose D2Snap – a first-of-its-kind downsampling algorithm for DOMs, intended for use as a DOM snapshot pre-processor. D2Snap does not rest on element extraction, but adopts an idea originated from signal processing: DOM nodes are locally consolidated, set to retain a majority of inherent UI features. Algorithmic output corresponds to a valid DOM itself.
We support D2Snap-downsampled DOM snapshots substantial performance: Downsampled snapshots are in the same estimated token size order of a grounded GUI snapshot baseline ($1\mathrm{e}{3}$). In respect to web-based tasks, the success rate of a GPT-4o backend providing interaction suggestions (67%) as well meets the baseline (65%). The best evaluated configuration – sized in $1\mathrm{e}{4}$, but within the model’s context window – outperforms the baseline by 8%. Our evaluation furthermore reveals that hierarchy represents a significant UI feature for LLMs. Image input – whether or not grounded – demonstrates little value for backend LLMs: performance of grounded GUI snapshots is close to grounding text alone.
2 DOM Snapshots
As highlighted, vast input token size constitutes the prevalent disadvantage of implementing DOM snapshots with web agents. On the contrary, five advantages stand out:
  1. 1.
    DOM serialises to HTML.
    LLMs are trained on vast amounts of HTML. Not least by agentic IDEs Anysphere Inc. ; Windsurf Inc , strong abilities to describe, classify, and navigate the inherent UI were supported Gur et al. (2022). Serialised DOM and HTML are structurally isomorphic MDN Contributors (c).
  2. 2.
    DOM virtualises for manipulation.
    Web agents are required to ground LLMs through (temporary) visual cues in the GUI. Collateral effects, such as flicker, apply He et al. (2024); Browser Use (a). By specification, DOMs can be cloned into memory MDN Contributors (h), for snapshots to be taken hidden from supervision.
  3. 3.
    DOM byte and token size correlate.
    In the majority of cases, screenshot data is several orders larger than DOM data Processing ; Fox Sports . Image input is pre-processed on LLM-side in order to disproportionate data and token size OpenAI (2025c). Either way, transfer of image data negatively affects round trip times.
  4. 4.
    DOM interaction is relative, not absolute.
    Vision capabilities alone constrain LLMs to target interaction by absolute means. If the layout shifts, action suggestions become obsolete. Pixel precision is not even granted yet OpenAI (2025c). DOMs allow programmatic targeting, such as with CSS selectors MDN Contributors (n).
  5. 5.
    DOM provides for early access.
    A screenshot is delayed by initial rendering of the GUI, which is signalled by the document load event MDN Contributors (d). By design, the DOM – a structural GUI requirement – is available beforehand, signalled by the document DOMContentLoaded event MDN Contributors (m).
3 DOM Downsampling
Rooted in signal processing, downsampling defines a technique for reducing data that scales out of time or space constraints. Chunks of data are thereby locally consolidated, while assuming relevant information is retained to a high degree O’Shea and Nash (2015). Broadly speaking, a JPEG image stores only an average colour for patches of pixels Wallace (1992). Effects of such downsampling visually increase with patch size, whereas the depicted object keeps being recognisable up to a large patch size [^3] The cover page conceptualises downsampling through an example image, and analogously for HTML (serialised DOM). . Related concepts are, in fact, implemented behind LLM vision APIs to subsidise image input OpenAI (2025c).
3.1 Downsampled DOM Snapshots
We herein propose D2Snap (Downsampled DOM Snapshot, or DOM to Snapshot): a first shot at downsampling applied to DOMs. Meant for use with LLM-based web agents, consolidation assumes to retain a majority of inherent UI features. By our definition, a UI feature is declarative information that perceptibly helps users solve tasks in scope of the respective application. Users, to that extent, comprise human and computer agents.
The DOM structurally resembles a tree. Each node represents a semantic entity. Both node syntax, and individual content may imply a UI feature – a button element with text “Submit” is likely clickable Haine (2007). There are three serialisable types of redundant DOM nodes: element, text, and attribute [^4] For consistency, our algorithmic design treats attributes as element node child nodes, and furthermore DOM leaf nodes.  MDN Contributors (l). We tie downsampling, i.e., consolidation, to node syntax. The degree to which a node represents a UI feature is not exclusively an objective matter. To reconnect with reported HTML interpretation strengths of LLMs, we draw ground truth about HTML (DOM) semantics from the latest GPT-4o (gpt-4o-2024-11-20OpenAI (c). The ground truth ultimately constitutes a UI feature degree rating per node type. Element nodes are, moreover, sub-classified by high-level purpose. While ratings could be perceived as algorithmic input, we herein imply those as constants. Attachment A lists the complete ground truth, including seminal model prompts.
INPUT: DOM; k, l, m $\in$ $[0,1]$
OUTPUT: DOM
PROCEDURE D2Snap $\in$ OO(|DOM|):
for NODE of post-order-Traversal(DOM):
switch Type(NODE):
case element’:
D2SnapElement(NODE, k)
break
case text’:
D2SnapText(NODE, l)
break
case attribute’:
D2SnapAttribute(NODE, m)
break
default:
Remove(NODE)
return DOM
Listing 1: Treversing the DOM tree, D2Snap handles nodes through a type-respective sub-procedure. A ratio of nodes and node contents is subtractively mutated in-place. The DOM is traversed in post-order to handle text nodes first, which prevents text downsampling from messing with higher level formatting. Parameters k, l, m influence the downsampling ratio per node type. The output corresponds to valid DOM itself.
D2Snap is a DOM traversal algorithm that applies a type-sensitive sub-procedure per node. We describe our algorithm through in-place mutations, since DOMs allow for being deep cloned into memory. To enable variable downsampling ratios, we introduce three ratio parameters on the unit interval, one for each sub-procedure. Listing 1 describes a high-level pseudocode view of D2Snap. Attachment B contains a comprehensive, recursive description of the algorithm.
3.2 Element Downsampling
Aligned with human commonsense, the ground truth tells apart three classes of UI feature elements: container, content, and interactive. Other elements (e.g., template MDN Contributors (k)) fall into a neglectable remainder class.
INPUT: ELEMENT, k $\in$ $[0,1]$
PROCEDURE D2SnapElement:
switch Class(ELEMENT):
case container’:
if Depth(ELEMENT) % k == 0
or Class(Parent(ELEMENT)) not container’:
break
E $\leftarrow$ { ELEMENT, Parent(ELEMENT) }
TARGET $\leftarrow$ $\operatorname*{argmin}_{e}$ { Semantics(e) | e $\in$ E }
SOURCE $\leftarrow$ CANDIDATES \ TARGET
Merge(SOURCE, TARGET)
break
case content’:
TEXT_NODE $\leftarrow$ TextNode(Markdown(ELEMENT))
replace(ELEMENT, TEXT_NODE)
break
case interactive’:
break
default:
Remove(ELEMENT)
Listing 2: D2Snap mutates element nodes based on their sub-classification: Container elements are merged depth-wise. Depth thereby corresponds to the total DOM height ratio, depicted by parameter k. According to ground truth, the name of the higher rated element is preserved. Attribute sets are joined, and collisions resolved in favour of the higher rated alternative. Content elements are translated to a less verbose Markdown representation. Interactive elements are kept as is. Other elements are considered noise, and hence removed.
Container
[…]
Element merge represents a binary operation: resolve naming collisions, and migrate child nodes. The element node whose type rates higher according to ground truth is the designated merge target. To retain valid HTML, only the target’s name is preserved. Attribute sets are unified, and collisions resolved in favour of the target. Child nodes are detached from the source, and reattached to the target node. For a bottom-up merge, this means, prepend the source node within the target node’s children. For a top-down merge – when the target is a child of the source node – prepend the target node with source children placed in front of it, and append the remaining children behind it. The target is subsequently reattached right in front of the source node. In any case, removal of the source node completes a merge.
Content
elements dictate a UI’s text formatting – not to be confused with individual text. Formatting induces nuanced semantics. Consider bold text, where stroke weight emphasises a word. Or a table, which spatially relates words among each other. D2Snap translates content elements to Markdown [^5] Markdown is a semantic equivalent to a the content element subset of HTML. We respect an extended Markdown flavour that includes, i.a., tables GitHub (2019).  Gruber and Swartz (2004). At that, we utilise its idea of being significantly more comprehensive than HTML. This works, as mixed code interpretation abilities of LLMs were previously supported Pian et al. (2023). Structurally, content translation corresponds to replacing the element by a single text node that contains its Markdown equivalent.
Interactive
elements represent – if not exclusively – a UI’s actuation interface. Interactive elements have dynamic handlers attached that are invoked when a user performs a certain action, such as a click. D2Snap retains all interactive elements as is, to allow LLMs direct suggestion of interaction targets. Notably, interactive elements with a Markdown representation are excluded from the above-sketched content downsampling strategy. For basic Markdown, this only holds true for hyperlink anchor elements (a). Listing 2 is a pseudocode description of the element node sub-procedure.
3.3 Text Downsampling
Text nodes contain actual natural language that renders in a UI. Downsampling of natural language boils down to eliminating units of text. Most common natural units of text are space-separated words, or punctuation-separated sentences. For D2Snap, we utilise the TextRank algorithm to rank sentences in a text node Mihalcea and Tarau (2004), and eliminate the least relevant sentences (lowest cumulative word entropy). A parameter l depicts the ratio of sentences to eliminate. For example, a text node with 5 sentences would result in 3 sentences for l = 0.5, or 1 sentence for l = 0.1. Listing 3 is a pseudocode description of the text node sub-procedure.
INPUT: TEXT, l $\in$ $[0,1]$
PROCEDURE D2SnapText:
TEXT_CONTENT $\leftarrow$ Content(TEXT)
SENTENCES $\leftarrow$ Tokenize(TEXT_CONTENT)
RANKED_SENTENCES $\leftarrow$ TextRanksentence(SENTENCES)
SELECTED_SENTENCES = Slice(
RANKED_SENTENCES,
[…]
Content(TEXT, Join(SELECTED_SENTENCES))
Listing 3: D2Snap splits contents of text nodes into a list of sentences. Sentences are ranked by relevance using the TextRank algorithm Mihalcea and Tarau (2004). The lowest ranking fraction of sentences is removed, decided by parameter l.
3.4 Attribute Downsampling
Not only element names, but also attribute names have UI semantics. This is obvious for attributes like disable, and inversely crossorigin MDN Contributors (i, j). Based on the ground truth, D2Snap filters attributes that score above a given threshold denoted by a parameter m. Unrated attributes, such as generic data attributes or artificial attributes, are considered to have zero semantics. Listing 4 is a pseudocode description of the attribute node sub-procedure. Attachment C contains a serialised DOM instance, followed by differently configured D2Snap downsampling results.
INPUT: ATTRIBUTE, m $\in$ $[0,1]$
PROCEDURE D2SnapAttribute:
if Semantics(ATTRIBUTE) < m:
Remove(ATTRIBUTE)
Listing 4: D2Snap removes attributes that, in the ground truth, rate below a threshold, denoted by parameter m. Attributes are identified by name.
3.5 Adaptive D2Snap
No parametric configuration of D2Snap guarantees a downsampled DOM below a specific size limit. We define, as a proof-of-concept, an algorithmic D2Snap-wrapper to downsample with adaptive ratio. Based on an initial DOM-respective estimate, AdaptiveD2Snap iterates a cyclic alternation of progressively increasing parameters. Our cyclic alternation rests on Halton Sequences, granting low-discrepancy in low dimensions Halton (1960). Bases in a sequence should empirically be adjusted so as to grow faster on parameters bound to less impactful UI features. With each iteration, a fundamental DOM size factor grows super-linearly. The factor bases on one megabyte as a soft DOM size upper-bound. Listing 5 describes AdaptiveD2Snap with pseudocode.
INPUT: DOM; tmax, imax, $\in\mathbf{N}$
OUTPUT: DOM
PROCEDURE AdaptiveD2Snap:
M $\leftarrow$ 1,000,000
i $\leftarrow$ 1
s $\leftarrow$ |DOM|
while true:
MAGNITUDE $\leftarrow$ $\frac{s}{M}$
HALTON_POINT $\leftarrow$ HaltonSequence(i++)
k $\leftarrow$ min{MAGNITUDE HALTON_POINT1, 1}
l $\leftarrow$ min{MAGNITUDE HALTON_POINT2, 1}
m $\leftarrow$ min{MAGNITUDE HALTON_POINT3, 1}
DOM_SNAPSHOT $\leftarrow$ D2Snap(DOM, k, l, m)
s $\leftarrow$ s1.125
if |DOM_SNAPSHOT| $\leq$ tmax:
return DOM_SNAPSHOT
if i $\geq$ imax:
error
Listing 5: AdaptiveD2Snap is an algorithmic wrapper for D2Snap. As long as the downsampled DOM is above a given input token size threshold tmax, the algorithm cycles progressively increasing parameter configurations k, l, m. A fundamental DOM size that influences the magnitude is therefore magnified per iteration. Cyclic alternation is built on a low-discrepancy Halton Sequence Halton (1960). An iteration count limit of imax ensures termination.
4 Evaluation
Generalist web agent evaluations do at most allow vague assumptions about the underlying snapshot utility. Agent specifics, not limited to system prompts, bear an undeniable threat to validity. However, existing datasets base on common web-based tasks He et al. (2024); Zhou et al. (2024); Xue et al. (2025); Deng et al. (2023), which favour abstraction of a dataset for isolated snapshot evaluation.
4.1 Dataset
Besides web tasks, a snapshot evaluation dataset needs to provide web application state across solution trajectories of distinct UIs – working around iterative agent logic. From the Online-Mind2Web dataset Xue et al. (2025); Deng et al. (2023), we randomly pick 6 easy, 6 medium, and 6 hard tasks.
We ask a third individual to solve each task under supervision. For each notably different UI, we serialise the respective web application state – both GUI, and DOM. We also serialise a grounded GUI snapshot close to how it is done by Browser Use [^6] We capture grounded GUI snapshots as follows: Visible interactive elements are enhanced with a coloured bounding box and a numerical identifier. The identifier is listed as text, supplemented with element tag name and contained text. We source our script from the Browser Use repository Browser Use (b) (https://github.com/surfly/D2Snap/blob/3788eb5d6f7d056d4a1f22cd100f0eea79d7fc27/snapshots/_highlight.js)  Müller and Žunič (2024). As an example, Attachment D displays a (grounded) GUI snapshot from the evaluation dataset. The seed dataset expands to a total of 52 records. A record, in other words, represents a partial solution space regarding a comprehensive web task.
In the next step, we let two individuals, both with a background in web development, independently annotate every record as follows: Identify all ways to (partially) solve the given task with the given snapshot. For each way, note down all alternative sets of elements that are required to interact with. For GUI snapshots, provide bounding box coordinates. For DOM snapshots, provide a unique CSS selector. For grounded snapshots, provide a numerical identifier (if available). Reference is given by joint annotations, which favourably substantiate agreement ($\kappa$ $\approx$ 0.7).
4.2 Procedure
To control the effects of different model system prompts, we create a short, generic template OpenAI (d). The system prompt template requests all elements that are required to interact with in order to solve an incidental web task. It is substituted with regard to snapshot class specifics: elements are supposed to be targeted in a format that aligns with the snapshot, such as CSS selectors. Attachment E shows the system prompt template, as well as concrete substitutes per snapshot class. Our evaluation framework iterates through the dataset, and prompts the latest GPT-4o (gpt-4o-2024-11-20OpenAI (b, c) with different subject snapshots, and the respective system prompt. If the set of suggested elements corresponds to a superset of any set in the record- and snapshot-respective reference, we count it a success [^7] During evaluation, elements are compared with slight tolerance: For GUI snapshots, we inflate the referenced bounding boxes by 10 pixels. For DOM-based snapshots, we as well accept a referenced parent, child or sibling node. .
4.3 Subjects
We consider grounded GUI snapshots, as implemented by Browser Use, the baseline snapshot technique. In our evaluation, we compare raw GUI snapshots, raw DOM snapshots (cut-off at 8,192 tokens), and parametrically diverse configurations of D2Snap-downsampled DOM snapshots:
  1. 1.
    GUI
  2. 2.
    DOM
  3. 3.
    GUIgrounded    Baseline
  4. 4.
    GUIgrounded   ||\ GUIgrounded $\setminus\ IMG$
  5. 5.
    D2Snapk(,l,m)   $|\ k,l,m\in[0,1]$
  6. 6.
    D2Snap     $|\ k\rightarrow\infty,l=0,\forall m$
  7. 7.
[…]
4.4 Results
Figure 1: Success rate per subject evaluated across the dataset. The rate of grounded GUI snapshots (GUIgr.) represents the baseline. Except for adaptive configurations, all D2Snap subjects outperform the baseline. Limited at five iterations, AdaptiveD2Snap is able to downsample about two thirds of DOMs in the dataset for a strict token limit of 8,192. With a limit of 32,768, it is able to downsample without error. Hierarchy supports highest utility across the assessed UI features.
In the first place, our evaluation results – averaged from three runs – support D2Snap as a useful pre-processor for DOM snapshots. D2Snap compares in success rate (67%) with the baseline (65%), and also in mean estimated token size (within $1\mathrm{e}{3}$). One token order higher, D2Snap.1, D2Snap.7, and D2Snap.6,.9,.3 (73%) significantly outperform the baseline by 8%. D2Snap.1 shows that low downsampling ratios already fit the entirety of DOMs from the dataset into the model’s context window (128K). In general, success rates with increasing parameters seem stable, but only drop towards asymptotes. On the other hand, mean token and byte input size decrease – both strongly correlate for DOM snapshots (r: 0.9994, P-value: $2.9\mathrm{e}{-6}$). Whereas mean byte size for grounded GUI snapshots is in $1\mathrm{e}{6}$, all D2Snap subjects are within $1\mathrm{e}{4}$ ($\sim$96% smaller). Table 1 presents the full evaluation results. Figure 1 plots success rates across subjects.
Success Errors Tokens Bytes
% % $\overline{X}$ $\overline{X}$
GUI 0 0 2,294 2,349,326
GUIgrounded 65 0 3,754 2,384,067
GUIgrounded 63 0 1,461 5,842
DOM 38 0 8,121 32,483
D2Snap.1 73 0 24,352 97,409
D2Snap.4 71 0 19,156 76,625
D2Snap.7 73 0 17,358 69,432
D2Snap 67 0 7,178 28,712
D2Snap.9,.3,.6 65 0 16,828 67,310
D2Snap.6,.9,.3 73 0 18,943 75,771
D2Snap.3,.6,.9 71 0 11,487 45,949
D2Snap4096 29 58 2,838 11,350
D2Snap8192 42 53 5,667 22,666
D2Snap32768 70 0 13,360 53,438
Table 1: Snapshot evaluation results cover success rate (Success %), error rate (Errors %), mean input token size (Tokens $\overline{X}$), and mean input byte size (Bytes $\overline{X}$). Grounded GUI snapshots (GUIgrounded) represent the baseline for our evaluation. Linearisation (D2Snap) achieves a mean DOM token size order equal to the baseline ($1\mathrm{e}{3}$), with slightly better performance. Our best configurations (i.a., D2Snap.6,.9,.3) outperform the baseline by a margin (8%). Limited at five iterations, D2Snap8192 fails to downsample roughly a third of DOMs from the dataset. D2Snap32768 downsamples without error. Within $1\mathrm{e}{5}$, all DOMs can be downsampled adaptively. Snapshots compiled from only grounding text (GUIgrounded) meet performance of grounded snapshots.
Figure 2: Comparison of mean input size across subjects, and estimated token and byte size per subject snapshots. Both token and byte size strongly correlate for text-modal D2Snap subjects. For the baseline, grounded GUI snapshots (GUIgr.), byte size scales way beyond remotely processed token size (hidden costs).
Figure 3: Estimated input token size per snapshot created by the subject D2Snap.6,.9,.3 across the entire dataset, sorted in ascending order. Greater part of snapshots meets the token order of the grounded GUI snapshots baseline. Only about a fifth of DOMs scale beyond an order of $1\mathrm{e}{4}$.
Adaptivity.
D2Snap8192 could downsample roughly two thirds of DOMs from the dataset below a strict token limit of 8,192. Fixed-configurations support that raising the input token limit renders adaptive downsampling highly reliable (below $1\mathrm{e}{5}$, i.e., the model context window). Figure 2 compares mean input token and byte sizes per, and across subjects. As illustrated in Figure 3, D2Snap.6,.9,.3 was able to output most snapshots within an order of $1\mathrm{e}{3}$ estimated input tokens. AdaptiveD2Snap indeed handles all DOMs from the dataset without error. D2Snap32768 outperforms the baseline by 5%.
Hierarchy.
Three evaluated parameter variations over the values 0.3, 0.6, and 0.9 give an idea of which assessed UI feature holds most value for web agent backend LLMs. Interestingly, hierarchy is revealed to be the strongest among those features; a high hierarchical downsampling ratio results in the lowest success rate among variations. D2Snap’s functional codomain arguably spans a discrete range between original HTML and fully linearised content – including interactive elements. Text-preserving linearisation compares to a reader view that covers all contents, not only contents within a main section.
Vision.
LLM vision capabilities, we observe, have low impact on snapshots. Success rates between GUIgrounded (65%), and GUIgrounded (snapshots compiled from only grounding text, i.e., without screenshots) (63%) are close to one another. Image data size, however, represents significant overhead.
5 Related Work
5.1 Element Extraction
Element extraction, with regard to web agents, characterises a technique to filter relevant parts from a DOM. It is usually integrated with grounded GUI snapshots. At that, extracted elements map to visual cues. Element extraction renders low-dimensional element arrays, disregarding hierarchy as a UI feature.
Müller and Žunič filter interactive elements by designated tag names. I.a., information retrieval tasks – draw data from any element – are out of such extraction scope. He et al. deploy an LLM to suggest relevant elements for a given task. Deng et al. let an LLM rank elements to select the top-k. Kim et al. implement a pre-trained model critique loop to improve rankings. Gur et al. rank elements by serialisation weight. Additional inference steps negatively affect round trip times. Extraction is an attempt to weakly solve the underlying agent problem, already. Sridhar et al. define extraction based on a Bayesian predictor. Element relevancy, to that extent, depends on previously performed actions. This approach moves towards a general web interaction model, but reportedly performs with low success.
5.2 Accessible Representations
Reader views and accessibility trees have existed to enhance readability of web application UIs for humans. Such representations are retrieved directly from a DOM, but no longer resemble HTML, which disposes of inherent LLM UI interpretation capabilities.
Reader views linearise main content of web pages Mozilla ; Ango . By design, reader views eliminate elements aside of a considered main content section. Besides, reader views flatten hierarchy in the fundamental DOM. Accessibility trees keep hierarchy intact, and, in fact, serve a purpose in line with LLM web agent backends World Wide Web Consortium (2023a, b); MDN Contributors (g). Detailed accessibility trees, however, transfer the DOM size problem to a higher abstraction layer.
6 Conclusion
DOM snapshots – if not for vast input size – promise several advantages over GUI snapshots. In this paper, we applied the concept of downsampling to DOMs. Ultimately, to enable DOM snapshots for web agents. Our approach notably differs from element extraction.
6.1 Contributions
We proposed D2Snap – an algorithm to downsample a DOM based on UI features. We support D2Snap substantial performance as a web agent snapshot pre-processor of DOMs. Our algorithm is able to downsize DOMs to an input token order of equivalent GUI snapshots ($1\mathrm{e}{3}$). Based on such D2Snap-downsampled DOM snapshots, an LLM (GPT-4o) is able to provide interaction suggestions for a diverse set of web tasks. The success rate (67%) compares to those of a grounded GUI snapshots baseline (65%). Our best evaluated parametric configuration (73%) outperforms the baseline by 8%. Either way, we showed that algorithmic pre-processing enables DOM snapshots for use with web agents. Moreover, we supported that hierarchy, which is retained with downsampled DOMs, represents a strong UI feature for LLMs. Element extraction techniques, in contrast, barely capture hierarchy. Lastly, success of only grounding snapshots (63%), i.e., disposing of screenshots, hints that images perform poorly as snapshots.
The repository related to this paper is available at https://github.com/webfuse-com/D2Snap [^8] The paper repository contains a D2Snap implementation, the evaluation dataset, and the evaluation framework (including a system prompt template)., [^9] The paper repository is redundantly archived at https://github.com/t-ski/D2Snap..
6.2 Limitations
Evaluation.
We crafted an evaluation dataset pivot to kick off snapshot isolated evaluations. To mitigate selection bias, we randomly took an even-difficulty subset of tasks from Online-Mind2Web. Our dataset, however, represents a relatively low sample size.
Application.
DOMs may embed other fully qualified DOMs. Web browsers render embeds, but do not grant programmatic access to cross-origin subtrees for security reasons World Wide Web Consortium . While GUI snapshots capture all of the rendered UI, DOM snapshots must not serialise beyond cross-origin subtree roots.
6.3 Future Work
Grounded GUI snapshots represent image-heavy, hybrid modality snapshots. Inversely, to evade certain limitations, (D2Snap-downsampled) DOM snapshots could conditionally be augmented with (scoped) screenshots to further elevate performance – from a substantial to a strong level.
Appendix A LLM Ground Truth
For D2Snap, we prompted GPT-4o OpenAI (b, c) to rate DOM concepts regarding their UI feature degree. The results represent universal downsampling ground truth.
A.1 Element Classification
Which classes of HTML tags exist with regard to semantics about the inherent user interface? Elements that have no semantics should be classified as ’other’.
A.2 Element Rating
For each existing HTML tag: Classify its type (’T’) based on a nominal scale with the categories ’interactive’, ’container’, ’content’, and ’other’. Score its semantics (’S’) with regard to how important the tag is for users to understand the inherent user interface, based on a rational scale on the unit interval [0, 1] for S. Respond with both the definition of each scale, and a list of triples (tag, T, S), ordered by T.
(’article’, container’, 0.95)
(’aside’, container’, 0.85)
(’body’, container’, 0.90)
(’div’, container’, 0.30)
(’footer’, container’, 0.70)
(’header’, container’, 0.75)
(’main’, container’, 0.85)
(’nav’, container’, 0.80)
(’section’, container’, 0.90)
(’a’, interactive’, 0.85)
(’button’, interactive’, 0.80)
(’details’, interactive’, 0.60)
(’form’, interactive’, 0.75)
(’input’, interactive’, 0.70)
(’label’, interactive’, 0.50)
(’select’, interactive’, 0.65)
(’summary’, interactive’, 0.55)
(’textarea’, interactive’, 0.65)
(’address’, content’, 0.60)
(’b’, content’, 0.40)
(’blockquote’, content’, 0.65)
(’code’, content’, 0.60)
(’em’, content’, 0.50)
(’figure’, content’, 0.50)
(’figcaption’, content’, 0.45)
(’h1’, content’, 1.00)
(’h2’, content’, 0.95)
(’h3’, content’, 0.90)
(’h4’, content’, 0.85)
(’h5’, content’, 0.80)
(’h6’, content’, 0.75)
(’hr’, content’, 0.20)
(’img’, content’, 0.60)
(’li’, content’, 0.60)
(’ol’, content’, 0.55)
(’p’, content’, 0.60)
(’pre’, content’, 0.55)
(’small’, content’, 0.30)
(’span’, content’, 0.20)
(’strong’, content’, 0.50)
(’sub’, content’, 0.25)
(’sup’, content’, 0.25)
(’table’, content’, 0.70)
(’tbody’, content’, 0.65)
(’td’, content’, 0.50)
(’th’, content’, 0.65)
(’tr’, content’, 0.50)
(’ul’, content’, 0.55)
(’base’, other’, 0.10)
(’br’, other’, 0.05)
(’canvas’, other’, 0.20)
(’head’, other’, 0.10)
(’html’, other’, 0.10)
(’link’, other’, 0.05)
(’meta’, other’, 0.00)
(’noscript’, other’, 0.05)
(’script’, other’, 0.00)
(’source’, other’, 0.05)
(’style’, other’, 0.00)
(’template’, other’, 0.00)
(’title’, other’, 0.40)
(’track’, other’, 0.05)
(’video’, other’, 0.50)
(*, other’, 0.00) [ed.]
A.3 Attribute Rating
For each existing HTML attribute: Score its semantics (’S’) with regard to how important the tag is for users to understand the inherent user interface, based on a rational scale on the unit interval [0, 1]. Respond with a list of tuples (attribute, S).
(’alt’, 0.9)
(’href’, 0.9)
(’src’, 0.8)
(’id’, 0.8)
(’class’, 0.7)
(’title’, 0.6)
(’lang’, 0.6)
(’role’, 0.6)
(’aria-*’, 0.6)
(’placeholder’, 0.5)
(’label’, 0.5)
(’for’, 0.5)
(’value’, 0.5)
(’checked’, 0.5)
(’disabled’, 0.5)
(’readonly’, 0.5)
(’required’, 0.5)
(’maxlength’, 0.5)
(’minlength’, 0.5)
(’pattern’, 0.5)
(’step’, 0.5)
(’min’, 0.5)
(’max’, 0.5)
(’accept’, 0.4)
(’accept-charset’, 0.4)
(’action’, 0.4)
(’method’, 0.4)
(’enctype’, 0.4)
(’target’, 0.4)
(’rel’, 0.4)
(’media’, 0.4)
(’sizes’, 0.4)
(’srcset’, 0.4)
(’preload’, 0.4)
(’autoplay’, 0.4)
(’controls’, 0.4)
(’loop’, 0.4)
(’muted’, 0.4)
(’poster’, 0.4)
(’autofocus’, 0.3)
(’autocomplete’, 0.3)
(’autocapitalize’, 0.3)
(’spellcheck’, 0.3)
(’contenteditable’, 0.3)
(’draggable’, 0.3)
(’dropzone’, 0.3)
(’tabindex’, 0.3)
(’accesskey’, 0.3)
(’cite’, 0.3)
(’datetime’, 0.3)
(’coords’, 0.3)
(’shape’, 0.3)
(’usemap’, 0.3)
(’ismap’, 0.3)
(’download’, 0.3)
(’ping’, 0.3)
(’hreflang’, 0.3)
(’type’, 0.3)
(’name’, 0.3)
(’form’, 0.3)
(’novalidate’, 0.2)
(’multiple’, 0.2)
(’selected’, 0.2)
(’size’, 0.2)
(’wrap’, 0.2)
(’hidden’, 0.1)
(’style’, 0.1)
(’content’, 0.1)
(’http-equiv’, 0.1)
(*, 0.0) [ed.]
Appendix B D2Snap
Below is a recursive, in-place description of the D2Snap algorithm. Type-specific sub-procedures are contained in case statements. The algorithm recurs first in order to follow a post-order traversal.
[…]
Appendix C Downsampling Examples
Herein, an exemplary DOM serialisation is followed by D2Snap-downsampling results based on different parametric configurations. Each parameter depicts a downsampling ratio $($$\in[0,1]$); for hierarchy (k), text (l), and attributes (m). Ratio and resulting DOM size are anti-proportionally related.
<section class="container" tabindex="3" required="true" type="example"> <div class="mx-auto" data-topic="products" required="false"> <h1>Our Pizza</h1> <div> <div class="shadow-lg"> <h2>Margherita</h2> <p> A simple classic: mozzarela, tomatoes and basil. An everyday choice! </p> <button type="button">Add</button> </div> <div class="shadow-lg"> <h2>Capricciosa</h2> <p> A rich taste: mozzarella, ham, mushrooms, artichokes, and olives. A true favourite! </p> <button type="button">Add</button> </div> </div> </div> </section>
C.1 k=.3, l=.3, m=.3 (55%)
<section tabindex="3" type="example" class="container" required="true"> # Our Pizza <div class="shadow-lg"> ## Margherita A simple classic: mozzarela, tomatoes, and basil. <button type="button">Add</button> ## Capricciosa A rich taste: mozzarella, ham, mushrooms, artichokes, and olives. <button type="button">Add</button> </div> </section>
C.2 k=.4, l=.6, m=.8 (27%)
<section> # Our Pizza <div> ## Margherita A simple classic: <button>Add</button> ## Capricciosa A rich taste: <button>Add</button> </div> </section>
[…]
# Our Pizza ## Margherita A simple classic: mozzarela, tomatoes, and basil. An everyday choice! <button>Add</button> ## Capricciosa A rich taste: mozzarella, ham, mushrooms, artichokes, and olives. A true favourite! <button>Add</button>
Appendix D GUI Snapshot Example
Left: (Full-page) GUI snapshot. A GUI snapshot corresponds to a screenshot, possibly limited by the viewport (as seen by a human).
Right: Grounded GUI snapshot—image. The coloured bounding boxes represent visual cues, identified by numerical indices.
Next Page: Grounded GUI snapshot—text. The leading identifiers, associated with the visual cues, describe interactive elements through (tag) name and text contents.
[0] A ""
[1] BUTTON "Menu"
[2] BUTTON "Search GOV.UK"
[3] LABEL "Search"
[4] INPUT ""
[5] BUTTON "Search GOV.UK"
[6] A "Get access to your eVisa"
[7] A "Childcare account: sign in"
[8] A "Benefits"
[9] A "Births, deaths, marriages and care"
[10] A "Business and self-employed"
[11] A "Childcare and parenting"
[12] A "Citizenship and living in the UK"
[13] A "Crime, justice and the law"
[14] A "Disabled people"
[15] A "Driving and transport"
[16] A "Education and learning"
[17] A "Employing people"
[18] A "Environment and countryside"
[19] A "Housing and local services"
[20] A "Money and tax"
[21] A "Passports, travel and living abroad"
[22] A "Visas and immigration"
[23] A "Working, jobs and pensions"
[24] A "Changes to GOV.UK"
[25] A "Find a job"
[26] A "National Insurance"
[27] A "Departments"
[28] A "News"
[29] A "Guidance and regulation"
[30] A "Research and statistics"
[31] A "Policy papers and consultations"
[32] A "Transparency documents"
[33] A "HMRC services: sign in"
[34] A "Check MOT history of a vehicle"
[35] A "Tax your vehicle"
[36] A "Universal Credit"
[37] A "Foreign travel advice"
[38] A "Check your State Pension age"
[39] A "Childcare account: sign in"
[40] A "Student finance: sign in"
[41] A "Self Assessment tax returns"
[42] A "Apply for a passport"
[43] BUTTON "Yes this page is useful"
[44] BUTTON "No this page is not useful"
[45] BUTTON "Report a problem with this page"
[46] A "Benefits"
[47] A "Births, death, marriages and care"
[48] A "Business and self-employed"
[49] A "Childcare and parenting"
[50] A "Citizenship and living in the UK"
[51] A "Crime, justice and the law"
[52] A "Disabled people"
[53] A "Driving and transport"
[54] A "Education and learning"
[55] A "Employing people"
[56] A "Environment and countryside"
[57] A "Housing and local services"
[58] A "Money and tax"
[59] A "Passports, travel and living abroad"
[60] A "Visas and immigration"
[61] A "Working, jobs and pensions"
[62] A "Departments"
[63] A "News"
[64] A "Guidance and regulation"
[65] A "Research and statistics"
[66] A "Policy papers and consultations"
[67] A "Transparency"
[68] A "How government works"
[69] A "Get involved"
[70] A "Help"
[71] A "Privacy"
[72] A "Cookies"
[73] A "Accessibility statement"
[74] A "Contact"
[75] A "Terms and conditions"
[76] A "Rhestr o Wasanaethau Cymraeg"
[77] A "Government Digital Service"
[78] A "Open Government Licence v3.0"
[…]
Appendix E Evaluation System Prompt
To maximise a control variable role of the system prompt used in our evaluation, we define a generic template. The template contains variable tags ({{ <TAG> }}) for each snapshot specific part of prompt. Below are the template, and specific substitutes per snapshot class.
E.1 Template
# Identity You are an AI agent that solves web-based tasks on behalf of a human user. Besides the task, the user provides some serialised representation of the considered web applications state. > Assume that today is July 16, 2025. # Instructions The user provides you with a web-based task, and serialsied state of the web application (referred toas a snapshot) to solve the task with. A task may be iterative, so it may not be possible to solve the taks completely, but only partially with the given state. Based on the state representation, your goal is to suggest all elements required to interact with in order to solve the task. It is important that the list of elements corresponds to a complete interaction trajectory. High precision when referencing target elements is key in order to be able to reproduce the interactions on the respective user interface. ## Input ### Task The web-based task is denoted with the prefix TASK:‘, e.g. "TASK: Show 4-star hotels in Amsterdam". ### Snapshot {{ SNAPSHOT_DESCRIPTION }} # Output Follow these rules when considering an element for interaction: - In case there are multiple trajectories to solve the task, rank them in memory according to human-readibility and choose the highest ranked alternative - If there are alternative elements per trajectory which seem to do the same thing, choose the most expressive alternative - Suppose there are only point and click actions, so never imply any other interaction ## Schema {{ SCHEMA_DESCRIPTION }} # Examples Consider the web-based task "TASK: Calculate the sum of 2 and 3.". <user_query> TASK: Calculate the sum of 2 and 3. </user_query> <user_query> {{ EXAMPLE_SNAPSHOT }} </user_query> <assistant_response> {{ EXAMPLE_RESPONSE }} </assistant_response>
E.2 GUI
{{ SNAPSHOT_DESCRIPTION }}
You are provided with a screenshot, namely the rendered GUI.
{{ SCHEMA_DESCRIPTION }}
Target elements by their spatial center pixel coordinates. This means, refer to them through an x (horizontal) and a y (vertical) pixel coordinate relative to the origin, which is in the top left corner of the image.
‘‘‘ base64 data:image/png;base64,iVBORw0KGgoAAAANSUhEUgAAAMgAAACWCAMAAACs... ‘‘‘
{{ EXAMPLE_RESPONSE }}
‘‘‘ json [ { "elementDescription": "Field that contains the mathematical expression to be solved.", "x": 100, "y": 47 }, { "elementDescription": "Button that triggers the calculation of the provided mathematical expression.", "x": 100, "y": 197 } ] ‘‘‘
E.3 GUIgrounded
{{ SNAPSHOT_DESCRIPTION }}
You are provided with two means of input: 1. A screenshot of the browser with bounding boxes and related numeric identifiers. 2. A list of interactive elements with format ‘[index] type "text"‘. index is the numeric identifier, type is an HTML element type (button, input, etc.), and text is the element description. > Numeric identifiers across means of input are consistent.
{{ SCHEMA_DESCRIPTION }}
Target elements by their numeric identifiers as given across both means of input.
‘‘‘ base64 data:image/png;base64,iVBORw0KGgoAAAANSUhEUgAAAMgAAACWCAIAAAAU... ‘‘‘ ‘‘‘ html [0] INPUT "Type expression" [1] BUTTON "Solve" [2] A "" ‘‘‘
{{ EXAMPLE_RESPONSE }}
‘‘‘ json [ { "elementDescription": "Field that contains the mathematical expression to be solved.", "identifier": 0 }, { "elementDescription": "Button that triggers the calculation of the provided mathematical expression.", "identifier": 1 } ] ‘‘‘
E.4 DOM/D2Snap
{{ SNAPSHOT_DESCRIPTION }}
You are provided with HTML, namely a serialised DOM.
{{ SCHEMA_DESCRIPTION }}
Target elements by shortest unique CSS selector. If the element has an assigned data-uid attribute, respond with only the respective data attribute selector, e.g. ‘[data-uid="21"]‘.
{{ EXAMPLE_SNAPSHOT }}
‘‘‘ html <main> <h1>Calculator</h1> <input id="expression" class="field" type="text" placeholder="3 * 4"> <button id="submit" type="button" data-uid="3">Solve</button> <div id="result"> <span></span> </div> ‘‘‘
{{ EXAMPLE_RESPONSE }}
‘‘‘ json [ { "elementDescription": "Field that contains the mathematical expression to be solved.", "cssSelector": "#expression" }, { "elementDescription": "Button that triggers the calculation of the provided mathematical expression.", "cssSelector": "[data-uid=\"3\"]" } ] ‘‘‘